Hi,
I can do this problem by sight; meaning I know what numbers will work by looking at it, I do not know how to do them if I couldn't do it by sight though.
The problem is:
x-2y=4
I did x=4 and y=0.
I know by looking at this that x could equal zero and y could equal a -4.
I do not know how to do the problem other than by figuring out in my head what 2 numbers would leave me with the answer of 4.
I hope you can help me. Thank you.


Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48516)

x-2y=4
----------
You have not said what the instructions are for
this equation. It has an infinite number of
"solutions".
Usually you are asked to put the equation in slope-
intercept form which is y=(1/2)x-2
Now if you let x=0 you get y=(1/2)(0)-2, so y=-2
But you can substitute any number for "x" and
then solve for "y".
Don't know if this helps you.
Cheers,
Stan H.

Answer by checkley71(8403)

X-2Y=4
X=2Y+4 SETTING X=0 WE GET
0=2Y+4
2Y=-4
Y=-4/2
Y=-2 WHEN X=0 THEN Y=-2
SETTING Y=5 WE GET
X-2*5=4
X-10=4
X=4+10
X=14 WHEN Y=5 THEN X=14